From one hospital, a small family rides their car and travels 25 km in a direction 30° south of west and then 30 km toward the north to a second hospital in search of a vacancy to accommodate a severe-case CoViD-19 relative patient. What is the separation distance between the two hospitals?


1
Expert's answer
2021-10-24T18:24:47-0400

"x=25\cdot\cos30°=21.65\ (m)"


"y=40-25\cdot\sin30°=27.5\ (m)"


"l=\sqrt{21.65^2+27.5^2}=35 (m)" . Answer